OpenStudy (anonymous):

At the local grocery store, cans of beans are arranged in rows. The number of cans in each row forms an arithmetic sequence. There are 105 cans in row 1, 96 cans in row 2, 87 cans in row 3, and so on. One row contains 42 cans. Which row is this?

OpenStudy (faiqraees):

OKay so form an equation a+(n-1)d = u1 105+(2-1)d = 96 FInd d using this Then substitute that d in the equation 105+(n-1)d = 42

OpenStudy (anonymous):

i did it the simple way

OpenStudy (anonymous):

it goes down by 9 everytime

OpenStudy (anonymous):

so its row 8

OpenStudy (faiqraees):

Good

OpenStudy (faiqraees):

But that method is the logical method and you should know it because it will be used in higher mathematics

OpenStudy (anonymous):

yes i know thank you.